OpenStudy (anthonyym):

The same frequency that causes a 0.25-m string to vibrate in its sixth harmonic also causes a 0.96-m pipe that is open at both ends to resonant in its second overtone. The speed of sound in air is 345 m/s. What is this common resonant frequency of the string and the pipe?

OpenStudy (anthonyym):

I have for the string: lambda = L/3 v = lambda*f f = (345m/s) / (0.25m/3) = 4140 Hz For the pipe: lambda = (2/3)L f = (345m/s) / [(2/3)*0.96m] = 539 Hz

OpenStudy (anthonyym):

Update: Am I not able to find the frequency for the string because I don't have the tension, linear density, and mass of the string by the formula \[v=\sqrt{\frac{ T }{ m/L }}\]

OpenStudy (fwizbang):

You are given that the two frequencies are the same, so you only need to find it using the pipe.

Parth (parthkohli):

Yeah, just calculate the frequency of the second overtone of the pipe and you're done.\[\frac{3v}{2L}\]